I need help with problems such as the one below, which is a practice problem. If I can have someone show me step by step instructions I'm sure I can complete the rest of my math assignment.
-3x+8= -4
Thank you!!!

Answer by mananth(16946)

-3x+8= -4
..
you need to bring all the constants on right side and terms on the left sides.
whatever operation you do on one side the same has to be done on the other side.
8 goes to the right side
add -8 to both sides of the equation
-3x+8 -8 = -4 -8
+8 -8 cancel off & (-4 -8 add up)
-3x = -12
-3 has to be separated from x
divide by -3 on both sides
-3x/-3= -12/-3
x= 4
cheers!
m.ananth@hotmail.ca